OpenStudy (anonymous):

1 question, am i right? ? 15v2 + 11v 11v2 + 15v 5v2 + 3v <------ my answer 3v2 + 5v

OpenStudy (unklerhaukus):

where is the question?

OpenStudy (anonymous):

8 sqrt v^2 - sqrt v^4 - sqrt 9v^2 + sqrt 16v^4

OpenStudy (phi):

do it term by term. what is \( 8 \sqrt{v^2} \)

OpenStudy (anonymous):

512 v ? im not good with dis stuff

OpenStudy (phi):

sqrt(v*v) is v v*v is often written as v^2 so sqrt(v^2) = v the 8 does not change. so what is \[ 8 \sqrt{v^2} \]

OpenStudy (anonymous):

oh i see now. the answer is d .

OpenStudy (phi):

what is \[ 8 \sqrt{v^2} \]